Construct a 90% confidence interval to estimate the population mean using the data below.

Xbar=24 S=3.4 n=22

What assumption needs to be made about this population?

The 90% confidence interval for the population is from a lower limit of— to an upper limit of—
Answer #1

Z_{\frac{\alpha }{2}}=1.645 for 90%

Z_{\frac{\alpha }{2}}\frac{s}{\sqrt{n}}=1.645\times \frac{3.4}{\sqrt{22}}

= \frac{5.593}{4.69}

= 1.19

Lower limit =\overline{x}-Z_{\frac{\alpha }{2}}\frac{s}{\sqrt{n}}

  =24-1.19

  =22.81

Upper limit =\overline{x}+Z_{\frac{\alpha }{2}}\frac{s}{\sqrt{n}}

  =24+1.19

  =25.19